OpenStudy (anonymous):

can someone help me solve this problem? x^32.1^3/4^3=1

OpenStudy (amistre64):

by solve you mean find "x" right?

OpenStudy (anonymous):

could you rewrite the problem with parantheses?

OpenStudy (amistre64):

x^32 = 64 if i see it right

OpenStudy (amistre64):

hard to tell with the way its written

OpenStudy (anonymous):

re write it

OpenStudy (anonymous):

(x^32.1^3)/(4^3)=1 yes solve for X

OpenStudy (amistre64):

\[x^{32.1^{3}}\]?

OpenStudy (anonymous):

-Bangs head on keyboard-

OpenStudy (anonymous):

Use the equation button

OpenStudy (amistre64):

(x^3) * (2.1^3) ? [x^(3.2)] [1^3] ? etc...

OpenStudy (anonymous):

well 4^3 is easy enough to solve, than multiply the 1 on the right by that number to reduce the fraction.

OpenStudy (anonymous):

its x^2 and rite next too it is 2.1^3 (idk how to use the equation button)

OpenStudy (anonymous):

x^32.1^3/4^3=1 1^3/4^3 0 1/64 x^32 must be the reciprocal of 1/64 Hence x^32 = 64 Hence x = 64 ^(1/32) = 1.38788635...

OpenStudy (amistre64):

i get: x =abt 2.62882

OpenStudy (amistre64):

i think an exact answer would maintain the radicals tho

OpenStudy (anonymous):

x^#/32 = 1

OpenStudy (amistre64):

i typoed that all over dint i

OpenStudy (anonymous):

It isnt an exact answer as the three dots imply but Amistre if you raise your result to the 32nd power it is quite a large number.

OpenStudy (amistre64):

x^3 = (4^3)/(2.1^3) x^3 = (4/2.1)^3 ; ^1/3 each side x = 4/2.1

OpenStudy (anonymous):

*x^3/32 =1 note i'm assuming the dot is multiplication and not a decimal

OpenStudy (anonymous):

\[x^3/32\]

OpenStudy (anonymous):

\[\frac{x^3\times2.1^3}{4^3}=1\] \[x^=\frac{4^3}{2.1^3}\] \[x=\frac{4}{2.1}\]

OpenStudy (amistre64):

\[\frac{x^3\ \ 2.1^3}{4^3}=1\]

OpenStudy (anonymous):

now your talking!

OpenStudy (amistre64):

\[\frac{{\color{red}{x^3}}\ {\color{blue}{2.1^3}}} {\color{green}{4^3}}=1\]

OpenStudy (anonymous):

yeesss! lol

OpenStudy (amistre64):

multiply thru by 4^3; and divide out the 2.1^3 x^3 = \(\cfrac{4^3}{2.1^3}\) ; now cuberoot each side x = \(\cfrac{4}{2.1}\)

OpenStudy (anonymous):

\[x = \sqrt[3]{32}\]

OpenStudy (anonymous):

amistre:the 2 is not cubed

OpenStudy (amistre64):

or x = 40/21 1.90476190476190476 ... ----------- 21 | 40.000000 21 ---- 190 189 ---- 10,0 84 ---- 160 147 ---- 130 126 ----- 40